# Readers Prefer Outputs of AI Trained on Copyrighted Books over Expert Human Writers

### v3

@@ 18 Ian McEwan (✓) 36 Percival (✓)Everett
Figure 15,16, 17 and 18 shows emulations of the original excerpt written by Colson Whitehead, Ottessa Moshfegh, Jonathan Franzen and Han Kang. There are 3 distinct MFA emulations per author (MFA1, MFA2, MFA3). For the In-context Prompting setup we prompt Claude 3.5 Sonnet (AI1), Gemini 1.5 Pro (AI2) and GPT-4o(AI3) to produce excerpts that can be pitted against MFA written excerpts. In finetuning setup we finetune GPT-4o and compare the fine-tuned output AI4 against MFA1, MFA2, MFA3. The output from fine-tuned GPT-4o (AI4) was generated using default hyperparameters.
Figure 15: In AI Condition 1(In-context prompt) we contrast MFA1/2/3 vs AI1/2/3. In AI Condition 2(Finetuning) we contrast MFA1/2/3 vs AI4
Figure 16: In AI Condition 1(In-context prompt) we contrast MFA1/2/3 vs AI1/2/3. In AI Condition 2(Finetuning) we contrast MFA1/2/3 vs AI4
Figure 17: In AI Condition 1(In-context prompt) we contrast MFA1/2/3 vs AI1/2/3. In AI Condition 2(Finetuning) we contrast MFA1/2/3 vs AI4
Figure 18: In AI Condition 1(In-context prompt) we contrast MFA1/2/3 vs AI1/2/3. In AI Condition 2(Finetuning) we contrast MFA1/2/3 vs AI4
@@
S2.5 Dissecting Preference Evaluations from Experts
Figure 20 and 21 shows the writing quality preference evaluation results from expert readers. Expert preferences are often grounded in solid reasoning and they often agree on similar snippets to support their reasoning. For instance look at uncooked spaghetti as an overwrought/awkward metaphor highlighted by both Expert Reader 1 and 3 while tarot card predicting his fate highlighted by Expert Reader 2 and 3. Grounding preferences in reasoning helps us uncover their mental models. It is also worth noticing that the preference often shifts towards AI once it is finetuned on authors entire oeuvre. Fine-tuning helps get rid of the officious disembodied robovoice that is characteristic of ChatGPT. Here experts praise the model for idiosyncratic narrative voice. By default due to post-training guardrails GPT-4o generates a rather safe and somewhat polite text in the in-context prompting setup. But what is particularly impressive here is when finetuned the vocabulary of the model generated text veers towards ribald and somewhat profane text that is characteristic of Tulathimutte’s voice (particularly Rejection from where the text is drawn). In Figure 22 we see that experts prefer the MFA emulation as more faithful to Junot Diaz’s voice. In fact the excerpt written by Gemini-1.5-Pro is oddly poetic and rife with purple prose and cliches which is nothing like Junot Diaz’s voice that is often marked by a dazzling hash of Spanish, English, slang, literary flourishes, and pure virginal dorkiness as noted by all three expert readers. However when finetuned on authors’ entire oeuvre (Figure 23) we see that the model learns these references and uses them in a rather wry and humorous manner (papi chulo, fifty pendejas, little pito off, las plagas) that is appreciated by expert readers leading them to prefer AI over MFAs.
Figure 20: MFA vs AI emulation of Tony Tulathimutte, read by 3 expert readers all of whom agree that MFA written excerpt is superior in terms of writing quality.
Figure 21: MFA vs AI emulation of Tony Tulathimutte, judged by 3 expert readers all of whom agree that Finetuned AI written excerpt is superior in terms of writing quality
Figure 22: MFA vs AI emulation of Junot Diaz, judged by 3 expert readers all of whom agree that MFA written excerpt is superior in terms of stylistic fidelity.
Figure 23: MFA vs AI emulation of Junot Diaz, judged by 3 expert readers all of whom agree that Finetuned AI written excerpt is superior in terms of stylistic fidelity
@@
Figure 24: Pangram AI likelihood scores vs GPTZero AI Likelihood scores for MFA, In-context and Finetuned
@@
  • In-context prompting: 2,400 judgments (1,200 per outcome); 4,800 long-format rows.
  • Fine-tuning: 1,440 judgments (720 per outcome); 2,880 long-format rows.
Key variables include writer type $W_{i}\in\{\text{Human},\text{GPT-4o},\text{Claude},\text{Gemini},\text{GPT-4o-FT}\}$, reader type $J_{j}\in\{\text{Expert},\text{Lay}\}$, and for H3, the Pangram AI-detection score $s_{i}\in[0,1]$. Throughout, Human and Expert serve as reference categories unless otherwise specified. Standard errors use CR2 with readers as the clustering unit; rows within a judgment pair are not independent, but in this design clustering by reader (the source of repeated measures) is the dominant dependence.
@@
We test our preregistered hypotheses using fixed-effects logistic regression with CR2 cluster-robust standard errors. The CR2 estimator provides improved finite-sample performance compared to standard cluster-robust estimators, particularly important given our relatively small number of readers (28 experts, 131 lay readers).
@@
$\text{logit}\,P(Y_{ijk}=1)=\alpha+\sum_{m\in\mathcal{M}}\beta_{m}\mathbf{1}[W_{i}=m]+\gamma\mathbf{1}[J_{j}=\text{Lay}]+\sum_{m\in\mathcal{M}}\phi_{m}\mathbf{1}[W_{i}=m]\cdot\mathbf{1}[J_{j}=\text{Lay}]$ (2)
where $\mathcal{M}=\{\text{GPT-4o},\text{Claude},\text{Gemini}\}$ and Human serves as the reference category. The interaction terms $\phi_{m}$ capture differential preferences between expert and lay readers. The preregistered H1 contrast tests whether humans outperform the average of baseline LLMs:
@@
where $\eta(w,g)$ denotes the linear predictor (fitted log-odds) for writer $w$ and reader group $g\in\{\text{Expert},\text{Lay}\}$. The odds ratio $\text{OR}^{\text{H1}}_{g}=\exp(\Delta^{\text{H1}}_{g})$ quantifies the relative preference, with values <1<1 indicating human preference and values >1>1 indicating AI preference.
@@
$\text{logit}\,P(Y_{ijk}=1)=\alpha+\beta_{\text{FT}}\mathbf{1}[W_{i}=\text{GPT-4o-FT}]+\gamma\mathbf{1}[J_{j}=\text{Lay}]+\phi_{\text{FT}}\mathbf{1}[W_{i}=\text{GPT-4o-FT}]\cdot\mathbf{1}[J_{j}=\text{Lay}]$ (4)
@@
All contrasts and confidence intervals use Wald (normal) inference with the CR2 covariance matrix. Specifically, 95% CIs are computed as estimate $\pm 1.96\cdot$ SE on the log-odds scale, then exponentiated. Holm correction is applied across the two reader-group contrasts (Expert, Lay) within each outcome (style, quality) and hypothesis (H1, H2).
@@
$\text{logit}\,P(Y_{ijk}=1)=$ $\alpha+\beta_{1}s_{i}+\beta_{2}\mathbf{1}[\text{Setting}_{i}=\text{FT}]+\beta_{3}\mathbf{1}[J_{j}=\text{Lay}]$
$+\beta_{12}s_{i}\cdot\mathbf{1}[\text{Setting}_{i}=\text{FT}]+\beta_{13}s_{i}\cdot\mathbf{1}[J_{j}=\text{Lay}]$
$+\beta_{23}\mathbf{1}[\text{Setting}_{i}=\text{FT}]\cdot\mathbf{1}[J_{j}=\text{Lay}]$
$+\beta_{123}s_{i}\cdot\mathbf{1}[\text{Setting}_{i}=\text{FT}]\cdot\mathbf{1}[J_{j}=\text{Lay}]$ (6)
@@
Inter-rater agreement quantifies consistency beyond chance using Fleiss’ kappa, appropriate here as each pair was evaluated by multiple readers (3 experts and 5 lay readers):
@@
where $\bar{P}$ represents mean observed agreement and $\bar{P}_{e}$ expected agreement by chance. Expert readers achieved moderate agreement ($\kappa=0.41-0.67$) while lay readers showed minimal agreement ($\kappa=0.07-0.22$), reflecting the subjective nature of literary evaluation.
@@
Unless noted otherwise, counts below refer to reader-level long-format observations entering the logistic models (two rows per judgment—one row per alternative in the pair). In the in-context prompting setting, 150 human–AI pairs were evaluated (evenly split across three baseline models: 50 Human vs GPT-4o, 50 Human vs Claude 3.5 Sonnet, 50 Human vs Gemini 1.5 Pro), yielding 1,200 judgments per outcome and 2,400 long-format rows. In the fine-tuned setting, 90 human–AI pairs yielded 720 judgments per outcome and 1,440 long-format rows. Human rows are three times more frequent than any single AI baseline in the in-context prompting setting because every pair contains a Human excerpt, while the 150 pairs are split evenly across the three baseline models.
@@
style In_Context Human Expert 450
style In_Context Human Lay 750
style In_Context GPT-4o_baseline Expert 150
style In_Context GPT-4o_baseline Lay 250
style In_Context Claude_baseline Expert 150
style In_Context Claude_baseline Lay 250
style In_Context Gemini_baseline Expert 150
style In_Context Gemini_baseline Lay 250
@@
style Fine_tuned Human Expert 270
style Fine_tuned Human Lay 450
style Fine_tuned GPT-4o_finetuned Expert 270
style Fine_tuned GPT-4o_finetuned Lay 450
@@
quality In_Context Human Expert 450
quality In_Context Human Lay 750
quality In_Context GPT-4o_baseline Expert 150
quality In_Context GPT-4o_baseline Lay 250
quality In_Context Claude_baseline Expert 150
quality In_Context Claude_baseline Lay 250
quality In_Context Gemini_baseline Expert 150
quality In_Context Gemini_baseline Lay 250
@@
quality Fine_tuned Human Expert 270
quality Fine_tuned Human Lay 450
quality Fine_tuned GPT-4o_finetuned Expert 270
quality Fine_tuned GPT-4o_finetuned Lay 450
@@
This section presents the numerical results from the logistic regression models specified in Section S4. All models use CR2 cluster-robust standard errors with readers as the clustering unit. The reference category throughout is Human × Expert. These tables provide the complete statistical foundation for Figure 2 in the main text.
@@
Table 8 reports the full coefficient tables for all four primary models. The dramatic sign reversal between in-context prompting and fine-tuned settings is immediately apparent: negative coefficients for AI writers in in-context prompting (indicating human preference) become positive in fine-tuning (indicating AI preference). The writer type $\times$ reader type interactions in the in-context prompting models reveal that lay readers are substantially more favorable to AI-generated text than experts, a difference that disappears after fine-tuning.
Table 8: GLM coefficients with CR2 robust standard errors for all primary models. Negative coefficients indicate preference for the reference category (Human), while positive coefficients indicate preference for AI. The sign reversal between in-context prompting and fine-tuned models represents the core finding.
@@
Style — In-context Prompting (n = 2,400)
(Intercept) 0.901 0.148 6.098 1.07e-09
writer_typeGPT-4o_baseline -1.655 0.402 -4.120 3.78e-05
@@
reader_typeLay -0.826 0.169 -4.877 1.08e-06
writer_typeGPT-4o_baseline:reader_typeLay 1.468 0.446 3.292 9.95e-04
writer_typeClaude_baseline:reader_typeLay 1.428 0.336 4.249 2.15e-05
writer_typeGemini_baseline:reader_typeLay 2.211 0.475 4.650 3.32e-06
Style — Fine-tuned (n = 1,440)
@@
reader_typeLay -0.008 0.186 -0.042 0.967
writer_typeGPT-4o_finetuned:reader_typeLay 0.015 0.372 0.042 0.967
Quality — In-context Prompting (n = 2,400)
@@
writer_typeClaude_baseline -1.246 0.368 -3.390 6.997e-04
@@
reader_typeLay -1.197 0.196 -6.124 9.12e-10
writer_typeGPT-4o_baseline:reader_typeLay 3.065 0.518 5.919 3.23e-09
writer_typeClaude_baseline:reader_typeLay 1.723 0.420 4.105 4.04e-05
writer_typeGemini_baseline:reader_typeLay 2.594 0.473 5.480 4.25e-08
Quality — Fine-tuned (n = 1,440)
(Intercept) -0.314 0.122 -2.571 0.0102
writer_typeGPT-4o_finetuned 0.627 0.244 2.571 0.0102
reader_typeLay -0.129 0.157 -0.821 0.412
writer_typeGPT-4o_finetuned:reader_typeLay 0.258 0.314 0.821 0.412
@@
Table 9 presents the preregistered contrasts testing H1 and H2, corresponding to panels A-B in Figure 2. The odds ratios reveal a striking reversal: expert readers’ 6-8 fold preference for human writing in in-context prompting conditions (OR = 0.16 for style, 0.13 for quality) transforms into an 8-fold preference for AI in fine-tuning (OR = 8.16 for style). Lay readers show less dramatic but directionally consistent shifts.
Table 9: Primary contrasts testing H1 (baseline LLMs vs Human) and H2 (fine-tuned vs Human). Holm correction applied within each outcome and hypothesis across reader types. The OR column shows odds ratios with values <1<1 favoring humans and >1>1 favoring AI.
@@
style H1 Expert -1.852 0.308 3.16e-09 1.26e-08 0.157 [0.085, 0.290]
style H1 Lay -0.150 0.166 0.365 0.365 0.861 [0.621, 1.193]
style H2 Expert 2.101 0.276 2.19e-14 2.19e-14 8.163 [4.693, 14.198]
style H2 Lay 2.116 0.241 1.76e-18 1.76e-18 8.290 [5.155, 13.333]
quality H1 Expert -2.021 0.405 1.16e-07 2.33e-07 0.133 [0.063, 0.280]
quality H1 Lay 0.441 0.180 0.014 0.019 1.554 [1.092, 2.212]
quality H2 Expert 0.626 0.248 0.012 0.016 1.873 [1.161, 3.021]
quality H2 Lay 0.886 0.197 7.08e-06 1.42e-05 2.424 [1.644, 3.574]
@@
Table 10 reports the predicted probabilities displayed in Figure 2 panels C-D. These probabilities, derived from the inverse-logit transformation of linear predictors, show the convergence of expert and lay preferences after fine-tuning. While experts strongly prefer humans over all baseline models in in-context prompting settings (probabilities 0.17-0.43), fine-tuning elevates AI preference to approximately 0.74 for both reader groups.
Table 10: Predicted probabilities of selecting AI excerpts by model and reader type, corresponding to Figure 2C-D. Values above 0.5 indicate AI preference. Note the convergence of expert and lay preferences in fine-tuned models.

### v4

@@ 18 Ian McEwan (✓) 36 Percival Everett (✓)
Figure 15,16, 17 and 18 shows emulations of the original excerpt written by Colson Whitehead, Ottessa Moshfegh, Jonathan Franzen and Han Kang. There are 3 distinct MFA emulations per author (MFA1, MFA2, MFA3). For the In-context Prompting setup we prompt Claude 3.5 Sonnet (AI1), Gemini 1.5 Pro (AI2) and GPT-4o(AI3) to produce excerpts that can be pitted against MFA written excerpts. In fine-tuning setup we fine-tune GPT-4o and compare the fine-tuned output AI4 against MFA1, MFA2, MFA3. The output from fine-tuned GPT-4o (AI4) was generated using default hyperparameters.
Figure 15: In AI Condition 1(In-context prompt) we contrast MFA1/2/3 vs AI1/2/3. In AI Condition 2(Fine-tuning) we contrast MFA1/2/3 vs AI4
Figure 16: In AI Condition 1(In-context prompt) we contrast MFA1/2/3 vs AI1/2/3. In AI Condition 2(Fine-tuning) we contrast MFA1/2/3 vs AI4
Figure 17: In AI Condition 1(In-context prompt) we contrast MFA1/2/3 vs AI1/2/3. In AI Condition 2(Fine-tuning) we contrast MFA1/2/3 vs AI4
Figure 18: In AI Condition 1(In-context prompt) we contrast MFA1/2/3 vs AI1/2/3. In AI Condition 2(Fine-tuning) we contrast MFA1/2/3 vs AI4
@@
S2.5 Dissecting Preference Evaluations from MFA-trained readers
Figure 20 and 21 shows the writing quality preference evaluation results from expert readers. Expert preferences are often grounded in solid reasoning and they often agree on similar snippets to support their reasoning. For instance look at uncooked spaghetti as an overwrought/awkward metaphor highlighted by both MFA-trained Reader 1 and 3 while tarot card predicting his fate highlighted by MFA-trained Reader 2 and 3. Grounding preferences in reasoning helps us uncover their mental models. It is also worth noticing that the preference often shifts towards AI once it is fine-tuned on authors entire oeuvre. Fine-tuning helps get rid of the officious disembodied robovoice that is characteristic of ChatGPT. Here experts praise the model for idiosyncratic narrative voice. By default due to post-training guardrails GPT-4o generates a rather safe and somewhat polite text in the in-context prompting setup. But what is particularly impressive here is when fine-tuned the vocabulary of the model generated text veers towards ribald and somewhat profane text that is characteristic of Tulathimutte’s voice (particularly Rejection from where the text is drawn). In Figure 22 we see that experts prefer the MFA emulation as more faithful to Junot Diaz’s voice. In fact the excerpt written by Gemini-1.5-Pro is oddly poetic and rife with purple prose and cliches which is nothing like Junot Diaz’s voice that is often marked by a dazzling hash of Spanish, English, slang, literary flourishes, and pure virginal dorkiness as noted by all three MFA-trained readers. However when fine-tuned on authors’ entire oeuvre (Figure 23) we see that the model learns these references and uses them in a rather wry and humorous manner (papi chulo, fifty pendejas, little pito off, las plagas) that is appreciated by MFA-trained readers leading them to prefer AI over MFAs.
Figure 20: MFA vs AI emulation of Tony Tulathimutte, read by 3 expert(MFA-trained) readers all of whom agree that MFA written excerpt is superior in terms of writing quality.
Figure 21: MFA vs AI emulation of Tony Tulathimutte, judged by 3 expert(MFA-trained) readers all of whom agree that Fine-tuned AI written excerpt is superior in terms of writing quality
Figure 22: MFA vs AI emulation of Junot Diaz, judged by 3 expert readers(MFA-trained) all of whom agree that MFA written excerpt is superior in terms of stylistic fidelity.
Figure 23: MFA vs AI emulation of Junot Diaz, judged by 3 expert readers(MFA-trained) all of whom agree that Fine-tuned AI written excerpt is superior in terms of stylistic fidelity
@@
Figure 24: Pangram AI likelihood scores vs GPTZero AI Likelihood scores for MFA, In-context and Fine-tuned
@@
  • In-context prompting: 6,600 judgments; 13,200 long-format rows.
  • Fine-tuning: 4,320 judgments; 8,640 long-format rows.
Key variables include writer type $W_{i}\in\{\text{Human},\text{GPT-4o},\text{Claude},\text{Gemini},\text{GPT-4o-FT}\}$, reader type $J_{j}\in\{\text{MFA-trained},\text{College-educated}\}$, and for H3, the Pangram AI-detection score $s_{i}\in[0,1]$. Throughout, Human and MFA-trained readers serve as reference categories unless otherwise specified. Standard errors use CR2 with readers as the clustering unit; rows within a judgment pair are not independent, but in this design clustering by reader (the source of repeated measures) is the dominant dependence.
@@
We test our preregistered hypotheses using fixed-effects logistic regression with CR2 cluster-robust standard errors. The CR2 estimator provides improved finite-sample performance compared to standard cluster-robust estimators, particularly important given our relatively small number of readers.
@@
$\text{logit}\,P(Y_{ijk}=1)$ $=\alpha+\sum_{m\in\mathcal{M}}\beta_{m}\mathbf{1}[W_{i}=m]+\gamma\mathbf{1}[J_{j}=\text{College-educated}]$ (2)
$\quad+\sum_{m\in\mathcal{M}}\phi_{m}\mathbf{1}[W_{i}=m]\cdot\mathbf{1}[J_{j}=\text{College-educated}]$
where $\mathcal{M}=\{\text{GPT-4o},\text{Claude},\text{Gemini}\}$ and Human serves as the reference category. The interaction terms $\phi_{m}$ capture differential preferences between MFA-trained and college-educated general readers. The preregistered H1 contrast tests whether humans outperform the average of baseline LLMs:
@@
where $\eta(w,g)$ denotes the linear predictor (fitted log-odds) for writer $w$ and reader group $g\in\{\text{MFA-trained},\text{College-educated}\}$. The odds ratio $\text{OR}^{\text{H1}}_{g}=\exp(\Delta^{\text{H1}}_{g})$ quantifies the relative preference, with values <1<1 indicating human preference and values >1>1 indicating AI preference.
@@
$\text{logit}\,P(Y_{ijk}=1)$ $=\alpha+\beta_{\text{FT}}\mathbf{1}[W_{i}=\text{GPT-4o-FT}]+\gamma\mathbf{1}[J_{j}=\text{College-educated}]$ (4)
$\quad+\phi_{\text{FT}}\mathbf{1}[W_{i}=\text{GPT-4o-FT}]\cdot\mathbf{1}[J_{j}=\text{College-educated}]$
@@
All contrasts and confidence intervals use Wald (normal) inference with the CR2 covariance matrix. Specifically, 95% CIs are computed as estimate $\pm 1.96\cdot$ SE on the log-odds scale, then exponentiated. Holm correction is applied across the two reader-group contrasts within each outcome (style, quality) and hypothesis (H1, H2).
@@
$\text{logit}\,P(Y_{ijk}=1)=$ $\alpha+\beta_{1}s_{i}+\beta_{2}\mathbf{1}[\text{Setting}_{i}=\text{FT}]+\beta_{3}\mathbf{1}[J_{j}=\text{College-educated}]$
$+\beta_{12}s_{i}\cdot\mathbf{1}[\text{Setting}_{i}=\text{FT}]+\beta_{13}s_{i}\cdot\mathbf{1}[J_{j}=\text{College-educated}]$
$+\beta_{23}\mathbf{1}[\text{Setting}_{i}=\text{FT}]\cdot\mathbf{1}[J_{j}=\text{College-educated}]$
$+\beta_{123}s_{i}\cdot\mathbf{1}[\text{Setting}_{i}=\text{FT}]\cdot\mathbf{1}[J_{j}=\text{College-educated}]$ (6)
@@
Inter-rater agreement quantifies consistency beyond chance using Fleiss’ kappa, appropriate here as each pair was evaluated by multiple readers (3 MFA-trained and 19 in-context/21 fine-tuning college-educated general readers):
@@
where $\bar{P}$ represents mean observed agreement and $\bar{P}_{e}$ expected agreement by chance. MFA-trained readers achieved moderate agreement ($\kappa=0.41-0.67$) while college-educated general readers showed minimal agreement ($\kappa=0.08–0.26$), reflecting the subjective nature of literary evaluation.
@@
Unless noted otherwise, counts below refer to reader-level long-format observations entering the logistic models (two rows per judgment—one row per alternative in the pair). In the in-context prompting setting, 150 human–AI pairs were evaluated (evenly split across three baseline models: 50 Human vs GPT-4o, 50 Human vs Claude 3.5 Sonnet, and 50 Human vs Gemini 1.5 Pro). Because each pair was evaluated by 3 MFA-trained readers and 19 college-educated general readers, this yields 3,300 judgments per outcome and 6,600 long-format rows. In the fine-tuned setting, 90 human–AI pairs were evaluated; with 3 MFA-trained readers and 21 college-educated general readers per pair, this yields 2,160 judgments per outcome and 4,320 long-format rows. Human rows are three times more frequent than any single AI baseline in the in-context prompting setting because every pair contains a Human excerpt, while the 150 pairs are split evenly across the three baseline models. In the fine-tuned setting, Human and GPT-4o fine-tuned rows occur equally often.
@@
style In_Context Human MFA-trained 450
style In_Context Human College-educated 2850
style In_Context GPT-4o_baseline MFA-trained 150
style In_Context GPT-4o_baseline College-educated 950
style In_Context Claude_baseline MFA-trained 150
style In_Context Claude_baseline College-educated 950
style In_Context Gemini_baseline MFA-trained 150
style In_Context Gemini_baseline College-educated 950
@@
style Fine_tuned Human MFA-trained 270
style Fine_tuned Human College-educated 1890
style Fine_tuned GPT-4o_finetuned MFA-trained 270
style Fine_tuned GPT-4o_finetuned College-educated 1890
@@
quality In_Context Human MFA-trained 450
quality In_Context Human College-educated 2850
quality In_Context GPT-4o_baseline MFA-trained 150
quality In_Context GPT-4o_baseline College-educated 950
quality In_Context Claude_baseline MFA-trained 150
quality In_Context Claude_baseline College-educated 950
quality In_Context Gemini_baseline MFA-trained 150
quality In_Context Gemini_baseline College-educated 950
@@
quality Fine_tuned Human MFA-trained 270
quality Fine_tuned Human College-educated 1890
quality Fine_tuned GPT-4o_finetuned MFA-trained 270
quality Fine_tuned GPT-4o_finetuned College-educated 1890
@@
This section presents the numerical results from the logistic regression models specified in Section S4. All models use CR2 cluster-robust standard errors with readers as the clustering unit. The reference category throughout is Human $\times$ MFA-trained. These tables provide the complete statistical foundation for Figure 2 in the main text.
@@
Table 8 reports the full coefficient tables for all four primary models. The dramatic sign reversal between in-context prompting and fine-tuned settings is immediately apparent: negative coefficients for AI writers in in-context prompting (indicating human preference) become positive in fine-tuning (indicating AI preference).
[…]
Table 8: GLM coefficients with CR2 robust standard errors for all primary models. Negative coefficients indicate preference for the reference category (Human), while positive coefficients indicate preference for AI. The sign reversal between in-context prompting and fine-tuned models represents the core finding. In fine-tuned models, the writer-type $\times$ reader-type interaction is significant, indicating that college-educated general readers favor AI by a wider margin than MFA-trained readers.
@@
Style — In-context Prompting (n = 6,600)
(Intercept) 0.901 0.148 6.098 1.08e-09
writer_typeGPT-4o_baseline -1.655 0.402 -4.121 3.78e-05
@@
reader_typeCollege-educated -0.929 0.153 -6.053 1.43e-09
writer_typeGPT-4o_baseline:reader_typeCollege-educated 1.784 0.414 4.307 1.66e-05
writer_typeClaude_baseline:reader_typeCollege-educated 1.583 0.292 5.412 6.22e-08
writer_typeGemini_baseline:reader_typeCollege-educated 2.357 0.438 5.384 7.27e-08
Style — Fine-tuned (n = 4,320)
@@
reader_typeCollege-educated -0.356 0.154 -2.308 0.021
writer_typeGPT-4o_finetuned:reader_typeCollege-educated 0.713 0.309 2.308 0.021
Quality — In-context Prompting (n = 6,600)
@@
writer_typeClaude_baseline -1.246 0.368 -3.390 7.00e-04
@@
reader_typeCollege-educated -1.275 0.179 -7.127 1.03e-12
writer_typeGPT-4o_baseline:reader_typeCollege-educated 3.197 0.484 6.601 4.08e-11
writer_typeClaude_baseline:reader_typeCollege-educated 1.935 0.381 5.073 3.91e-07
writer_typeGemini_baseline:reader_typeCollege-educated 2.716 0.433 6.268 3.65e-10
Quality — Fine-tuned (n = 4,320)
(Intercept) -0.314 0.122 -2.571 0.010
writer_typeGPT-4o_finetuned 0.627 0.244 2.571 0.010
reader_typeCollege-educated -0.531 0.133 -3.997 6.41e-05
writer_typeGPT-4o_finetuned:reader_typeCollege-educated 1.062 0.266 3.997 6.41e-05
@@
Table 9 presents the preregistered contrasts testing H1 and H2, corresponding to panels A–B in Figure 2. The odds ratios reveal a striking reversal: MFA-trained readers’ 6–8 fold preference for human writing in in-context prompting conditions (OR = 0.16 for style, 0.13 for quality) transforms into an 8-fold preference for AI in fine-tuning (OR = 8.16 for style). College-educated general readers show directionally consistent but substantially larger shifts, particularly after fine-tuning (OR = 16.65 for style, 5.42 for quality), indicating that this reader group favors AI-generated text by a wider margin than MFA-trained readers.
Table 9: Primary contrasts testing H1 (baseline LLMs vs Human) and H2 (fine-tuned vs Human). Holm correction applied within each outcome and hypothesis across reader types. The OR column shows odds ratios with values <1<1 favoring humans and >1>1 favoring AI. Note: The p and $p_{\mathrm{Holm}}$ columns report exact two-sided Wald p-values and their Holm-corrected counterparts. The main text reports Holm-corrected p-values using inequality bounds ($p<10^{-k}$), capped at $p<10^{-15}$ as a conservative floor reflecting the practical limits of the normal approximation.
@@
(log-odds) (log-odds) exp(Est.) (OR scale)
style H1 MFA-trained -1.852 0.314 3.65e-09 7.31e-09 0.157 [0.08, 0.29]
style H1 College-educated 0.056 0.083 0.501 0.501 1.058 [0.90, 1.25]
style H2 MFA-trained 2.100 0.282 1.04e-13 1.04e-13 8.163 [4.69, 14.20]
style H2 College-educated 2.813 0.125 4.91e-112 9.81e-112 16.652 [13.03, 21.28]
quality H1 MFA-trained -2.020 0.381 1.19e-07 1.19e-07 0.133 [0.06, 0.28]
quality H1 College-educated 0.597 0.087 6.00e-12 1.20e-11 1.816 [1.53, 2.15]
quality H2 MFA-trained 0.627 0.244 1.02e-02 1.02e-02 1.873 [1.16, 3.02]
quality H2 College-educated 1.690 0.105 4.63e-58 9.27e-58 5.417 [4.41, 6.66]
@@
Table 10 reports the predicted probabilities displayed in Figure 2 panels C–D. These probabilities, derived from the inverse-logit transformation of linear predictors, show that while MFA-trained readers strongly prefer humans over all baseline models in in-context prompting settings (probabilities 0.17–0.43), fine-tuning elevates AI preference to 0.74 for MFA-trained readers and 0.80 for college-educated general readers on stylistic fidelity, and to 0.58 versus 0.70 on writing quality. Unlike the in-context condition, both reader groups now favor fine-tuned AI, but college-educated general readers do so by a significantly wider margin.
Table 10: Predicted probabilities of selecting AI excerpts by model and reader type, corresponding to Figure 2C–D. Values above 0.5 indicate AI preference. After fine-tuning, both reader groups prefer AI, but college-educated general readers show higher predicted probabilities than MFA-trained readers.